OpenStudy (anonymous):

simplify expression: 2/x+5 - 3x/x+3

OpenStudy (anonymous):

(6-x)/[(x+5)(x+3)]

OpenStudy (anonymous):

how did u solve that

OpenStudy (anonymous):

\[\frac {2(x+3)-3x(x+5)}{(x+3)(x+5)}=\frac{-9x^2-13x+6}{(x+3)(x+5)}\]